quantity of 40% acid solution and another quantity of 10% acid solution must be mixed to form 15L of solution that 20% acid. how much of each must be mixed?
Answer by ewatrrr(24785)


Hi,
Let x represent the amount of 40% solution
then (15L -x) will be the amount of the 10% solution
.40x + .10(15L-x)= .20*15L
solving for x
.30x = .10*15L
x = 5L of the 40% solution and 10L of the 10% solution
checking our answer
.40*5 + .10 *10 = 2 + 1 = 3= .20*15